OpenStudy (anonymous):

Find the equation of a line that meets the given condition : perpendicular to 6x-2y=10,passes through (6,-3)

OpenStudy (texaschic101):

first thing we have to do is find the slope of the given equation. Do you know how to do that ?

OpenStudy (anonymous):

No

OpenStudy (texaschic101):

put the equation , 6x - 2y - 10, into y = mx + b form. Can you do that ?

OpenStudy (texaschic101):

oops...6x - 2y = 10

OpenStudy (nirmalnema):

slope = -a/b

OpenStudy (texaschic101):

put your equation in y = mx + b form to find the slope

OpenStudy (texaschic101):

do you need help in doing this ?

OpenStudy (nirmalnema):

slope is -6/-2 =3

OpenStudy (nirmalnema):

so the slope of the line perpendicular to it is -1/3

OpenStudy (texaschic101):

the asker needs to learn how to find the slope

OpenStudy (anonymous):

I want you to tell me how to find slope explain the steps then I can learn

OpenStudy (texaschic101):

how to put your equation in y = mx + b form... 6x - 2y = 10 (subtract 6x from both sides) -2y = -6x + 10 (now divide by -2) y = -6/-2y + 10/-2 y = 3/1x - 5 your slope in y = mx + b form is in the position of m. therefore, your slope in this equation is 3/1. Do you understand so far ?

OpenStudy (texaschic101):

typo...y = -6/-2x + 10/-2

OpenStudy (anonymous):

Yes I do thanks

OpenStudy (texaschic101):

now we are looking for a perpendicular line, therefore we need to find the negative reciprocal of the slope 3/1. To find the negative reciprocal , we " flip " the fraction and change the sign. Therefore, the perpendicular slope we need is -1/3. Understand so far ?

OpenStudy (texaschic101):

@shakur ...you still with me ?

OpenStudy (anonymous):

Yes I was trying to work on similar equation using the steps you gave me

OpenStudy (texaschic101):

now we need to find the equation of the perpendicular line. To do this we will use this formula : y - y1 = m(x - x1) slope(m) = -1/3 (6,-3) x1 = 6 and y1 = -3 now lets sub y - (-3) = -1/3(x - 6) y + 3 = -1/3x + 2 y = -1/3x + 2 - 3 y = -1/3x - 1 <-- your equation do you understand ? do you have any questions ? I will be more then happy to answer them.

OpenStudy (anonymous):

I am fine,I will solved it now and let you if i got it accoring to your steps

OpenStudy (texaschic101):

Just remember, when you are looking for a perpendicular line, you need the negative reciprocal of the slope given. Now if you were looking for a parallel line, the slope would be the same as the slope given.